Find the number of terms in an A.P, whose first term is 1, last term is 77 and the sum of terms is 507
Answer by math_tutor2020(3817)


n = number of terms

Sn = sum of first n terms of arithmetic progression (AP)
Sn = (n/2)*(first term + nth term)
507 = (n/2)*(1 + 77)
507*2 = 78n
1014 = 78n
n = 1014/78
n = 13

Answer: 13 terms